OpenStudy (anonymous):

Rationalize the denominator and simplify 8x/ ^5sqrt(27x^17y^13) Is the answer 8^5sqrt(9x^3y^2)/3x^4y^3)

OpenStudy (ash2326):

Is this your question? \[\frac{8x}{\sqrt[5]{27x^17y^13}}\]

OpenStudy (anonymous):

Yes

OpenStudy (ash2326):

Let's simplify it \[\frac{8x}{\sqrt[5]{27x^{17}y^{13}}}\] Let's write x and y 's powers as multiples of 5 \[\frac{8x}{\sqrt[5]{27x^{15}x^2y^{10}y^3}}\] We get now \[\frac{8x}{x^3y^2\sqrt[5]{27x^2y^3}}\] Now if we want to remove root from the denominator then we need to multiply \((\sqrt[5]{27x^2y^3})^4\) to both numerator and denominator We get \[\frac{8x \times (\sqrt[5]{27x^2y^3})^4}{x^3y^2(27x^2y^3)}\] We get \[\frac{8x \times (\sqrt[5]{27x^2y^3})^4}{27x^5y^5}\] Now we get \[\frac{8 \times (\sqrt[5]{27x^2y^3})^4}{27x^4y^5}\]